### v1

@@ 2.1 Change Detection Methods
  • TTP [7] enhances high-precision change detection in complex spatio-temporal remote sensing scenarios by integrating the general knowledge of foundational visual models into the change detection task. It leverages the potential of foundational models even with limited annotated change detection data. TTP overcomes domain shift issues during knowledge transfer and addresses the challenge of expressing both homogeneous and heterogeneous features in multi-temporal images. Specifically, TTP utilizes general segmentation knowledge based on the segment anything model (SAM) by introducing low-rank fine-tuning parameters into the SAM backbone, which mitigates spatial semantic domain shifts. Furthermore, TTP proposes a time-travel activation gate, enabling temporal features to permeate the pixel semantic space, thus enhancing the foundational model’s ability to understand both homogeneous and heterogeneous features of bi-temporal images. Lastly, an efficient multi-level change prediction head is designed to decode dense and high-level semantic change features within the foundational model. This method paves the way for more accurate and efficient change detection in remote sensing imagery.
@@ 3.1 Overall
Config Files. In Open-CD, the model, optimizer, dataset, data piprline, etc. are specified in config files. In general, if the user does not need to redesign the model internally, simply modifying the config file is enough.
@@ 5.3 Downstream validation for generative model
Although huge amount of remote sensing images are available through Google Earth etc., data annotation remains a challenge, which is time-consuming and costly. One trend is to synthesize simulated image-label pairs through conditional GAN, Diffusion models, etc [38, 29, 39]. and use these data to (pre-)train downstream perception models. In this case, Open-CD can be applied to validate the quality of the simulated change detection data. A typical study is ChangeAnywhere [29], which proposes a diffusion model to generate a simulated change detection dataset ChangeAnywhere-100K and uses Open-CD for subsequent validation.

### v2

@@ 2.1 Change Detection Methods
  • TTP [7] enhances high-precision change detection in complex spatio-temporal remote sensing scenarios by integrating the general knowledge of foundational visual models into the change detection task. It leverages the potential of foundational models even with limited annotated change detection data. TTP overcomes domain shift issues during knowledge transfer and addresses the challenge of expressing both homogeneous and heterogeneous features in multi-temporal images. Specifically, TTP utilizes general segmentation knowledge based on the segment anything model (SAM) by introducing low-rank fine-tuning parameters into the SAM backbone, which mitigates spatial semantic domain shifts. Furthermore, TTP proposes a time-travel activation gate, enabling temporal features to permeate the pixel semantic space, thus enhancing the foundational model’s ability to understand both homogeneous and heterogeneous features of bi-temporal images. Lastly, an efficient multi-level change prediction head is designed to decode dense and high-level semantic change features within the foundational model. This method paves the way for more accurate and efficient change detection in remote sensing image.
  • MTKD [20] is an innovative approach to enhance change detection in remote sensing image. Existing algorithms often struggle to consistently deliver satisfactory results across the diverse scenarios of natural and human-induced changes. MTKD is designed to address this issue comprehensively. The training process of the MTKD framework is comprised of three stages. Initially, an original model is trained on the complete training set. Subsequently, the training data is partitioned based on the change area ratio (CAR), enabling the optimization of distinct teacher models for each specific range of changes. Finally, the student model, initialized with the parameters from the original model, is trained under the supervision of these specialized teacher models, integrating the diverse strengths taught by each teacher. This methodology significantly enhances the student’s ability to manage a wide range of change scenarios. The MTKD framework empowers the student model to achieve superior detection accuracy without additional computational or time costs during inference, marking a significant advancement in change detection by ensuring robust performance across various network architectures.
@@ 3.1 Overall
Config Files. In Open-CD, the model, optimizer, dataset, data pipeline, etc. are specified in config files. In general, if the user does not need to redesign the model internally, simply modifying the config file is enough.
@@ 5.3 Downstream validation for generative model
Although huge amount of remote sensing images are available through Google Earth etc., data annotation remains a challenge, which is time-consuming and costly. One trend is to synthesize simulated image-label pairs through conditional GAN, Diffusion models, etc. [39, 30, 40] and use these data to (pre-)train downstream perception models. In this case, Open-CD can be applied to validate the quality of the simulated change detection data. A typical study is ChangeAnywhere [30], which proposes a diffusion model to generate a simulated change detection dataset ChangeAnywhere-100K and uses Open-CD for subsequent validation.